find the indicated z-score shown in the graph

HW Score: 38.33%, 11.5 of 30 Question Help 0.0244 0.0244 Z=? o z=? Check Answer Next 11:33 PM 1/12/2020
Answer #1

Shay? , 0.0244 0.0244 2:0 22? 1.96 - 2:1-1096 since - 10.02.44 +0.0244) = 1-0.0488 = 0,95190.95 so zo score at left their sid